# Efficient HDR Reconstruction from Real-World Raw Images

### v5

@@ Abstract
The widespread usage of high-definition screens on edge devices stimulates a strong demand for efficient high dynamic range (HDR) algorithms. However, many existing HDR methods either deliver unsatisfactory results or consume too much computational and memory resources, hindering their application to high-resolution images (usually with more than 12 megapixels) in practice. In addition, existing HDR dataset collection methods often are labor-intensive. In this work, in a new aspect, we discover an excellent opportunity for HDR reconstructing directly from raw images and investigating novel neural network structures that benefit the deployment of mobile devices. Our key insights are threefold: (1) we develop a lightweight-efficient HDR model, RepUNet, using the structural re-parameterization technique to achieve fast and robust HDR; (2) we design a new computational raw HDR data formation pipeline and construct a real-world raw HDR dataset, RealRaw-HDR; (3) we propose a plug-and-play motion alignment loss to mitigate motion ghosting under limited bandwidth conditions. Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU. While being highly efficient, our model also outperforms the state-of-the-art HDR methods in terms of PSNR, SSIM, and a color difference metric.
1 Introduction
@@
Most resource-constrained cameras exhibit low dynamic ranges (LDR), rendering them unable to capture the full range of brightness and color information in real-world scenes. Conversely, high dynamic range (HDR) imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images. Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users. As a result, there has been an increasing focus on fusion-based HDR imaging methods.
@@
Recent methods [35, 20, 4, 25, 34] based on convolution neural networks (CNNs) [17, 22] have made impressive progress in HDR reconstruction performance, thanks to their scalability and flexibility from constructing elementary building blocks like convolutional layers. However, superior performance is usually obtained at a cost of heavy computational burden [20, 34, 21]. Although this can be alleviated by elaborate network structures or dedicated computing engines (e.g., GPU and NPU), the hardware cost and power consumption still limit the deployment of existing deep HDR reconstruction networks. Specifically, the growing number of high-definition screens on edge devices (e.g., smartphones, security cameras, and televisions) calls for a practical HDR reconstruction solution.
@@
On the other hand, in the image processing pipeline, HDR reconstruction is widely used in the sRGB domain. Previous methods [35, 22, 36] exploit a set of sRGB images with different exposure levels to produce an HDR image, which has made rapid development in recent years. However, they tend to overlook three critical aspects. 1) Dataset Collection: Existing methods [3, 46] follow Kalantari et al. [14] to construct datasets. They first make the subject static and take three sets of images with different exposures, and then make the subject move twice to take dynamic images with different exposures. However, this process is labor intensive and difficult to acquire on a large scale. 2) ISP Processing Speed: When obtaining raw LDR images with different exposures, the ISP pipeline must be performed separately on each exposure. This incurs additional memory and computational overhead and leads to lower frame rates for HDR image output. 3) Reconstruction Quality: Raw images contain more delicate details of the original sensor signal that can be lost while processing sRGB images. The limitations of current HDR reconstruction methods highlight the need for further research and development in HDR imaging.
@@
In this paper, we propose an efficient scheme for HDR image reconstruction in the raw image domain. By analyzing the HDR image sensor system, we design a lightweight and efficient model for raw HDR reconstruction named RepUNet. RepUNet adopts reparameterization techniques and does not contain explicit computationally expensive alignment modules, such as optical flow [14], deformation convolution [4], or attention [20, 37], which are commonly used in existing deep learning-based HDR reconstruction methods [14, 20, 33, 24]. To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss, which encourages the network to focus on local motion patterns and alleviate misalignment between short- and long-exposure images. Consequently, our approach significantly reduces hardware costs and improves the real-time performance of HDR imaging systems.
@@
To further promote the HDR imaging system, we investigate the HDR sensor imaging principle. We observe that changing the Gain of the image sensor can have a similar effect as modifying the exposure time under noise-free conditions. Leveraging this insight, we design an automatic control imaging system that captures raw images with different exposures, based on a digital camera photoelectric signal conversion model. This automatic control operable system satisfies real-world scenes’ dynamic range requirements, making it a practical tool for generating high-quality HDR images. The resulting RealRaw-HDR dataset includes many LDR-HDR pairs for training and evaluation. By incorporating the unique characteristics of raw images into our approach, we can achieve superior HDR reconstruction results with increased efficiency and accuracy.
@@
    We investigate the structure re-parameterizable technique for the HDR task and propose a lightweight model, RepUNet, with Topological Convolution Block (TCB). TCB can be used to improve the HDR performance of any HDR model without introducing any extra burden for inference.
@@
Our contributions represent a significant step forward in raw HDR image reconstruction research, providing an effective and efficient solution for producing high-quality HDR images. Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU. While highly efficient, our model also outperforms the state-of-the-art HDR methods by a large margin in terms of PSNR, SSIM, and a color difference metric.
2 Related Work
2.1 HDR Imaging
@@
Recently, benefiting from the fast development of deep learning techniques, training deep neural networks for effective HDR reconstruction has become increasingly popular. Many methods apply deep neural networks [16, 24, 29, 13] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images. Kalantari et al. [14] proposed a CNN-based HDR approach that employs optical flow to align LDR sRGB images before network inference. Wu et al. [33] approached HDR imaging as an image translation problem without explicit motion alignment. Yan et al. [34] introduced spatial attention to achieve LDR image alignment. Liu et al. [20] presented an attention-guided deformable convolutional network for multi-frame HDR imaging. Prabhakar et al. [27] introduced an efficient method for generating HDR images using a bilateral guided up-sampler and exploring zero-learning for HDR reconstruction. Niu et al. [25] proposed a multi-frame HDR imaging method based on generative adversarial learning. Liu et al. [21] proposed a Transformer-based [26] HDR imaging method. These deep learning-based approaches [34, 20, 21] consistently elevated state-of-the-art performance. However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset. They overlook the large computational and storage resources the ISP pipeline requires to process bracketed exposure raw images. It also complicates the ISP system, making it challenging for resource-limited cameras to output high-quality video/images.
@@
To simplify the ISP system, another class of HDR reconstruction methods is based on raw image input. Google HDR+ produced the raw HDR image by aligning and merging a burst of raw frames with the same low exposure. Nevertheless, this approach requires a complex ISP system design and takes up a lot of DDR memory. Zou et al. [46] proposed reconstructing HDR images from a single raw image and collecting a raw/HDR paired dataset. However, this dataset is not suitable for real HDR sensors. Therefore, none of the existing methods can meet the requirements of real scenarios.
2.2 Low-level Raw Image Processing
@@
Due to the merits of raw data, raw-based image processing [18, 41] has made significant progress in recent years. The work in [40] first performs the demoireing task in the raw domain and then utilizes a pretrained ISP module to transform the result into the sRGB domain. Yang et al. [38] proposed a single-stage network empowered by feature domain adaptation to decouple the denoising and color mapping tasks in raw low-light enhancement. Zhang et al. [42] constructed a real-world super-resolution dataset by designing an optical zoom system and proposed a baseline network with a bilateral contextual loss. Qian et al. [28] solved the joint demosaicing, denoising, and super-resolution task with the raw input. Wang et al. [30] proposed a lightweight and efficient network for raw image denoising. Sharif et al. [1] proposed a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors. Wei et al. [31] investigated the low-light image denoising considering the sensor photoelectric properties. Yue et al. [39] achieved state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs. Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data. However, acquiring paired data in the raw domain is difficult and expensive. Our work proposes a new large-scale, high-quality raw dataset and provides a pipeline to acquire raw LDR-HDR pairs based on the imaging system.
3 New raw LDR-HDR pair Formation Pipeline
@@
3.1 Analysis of CMOS Imaging System
@@
Subsequently, the camera amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [45]:
@@
where $K_{\textrm{a}}$ represents the Analog Gain, $K_{\textrm{d}}$ stands for the Digital Gain, and $V_{\textrm{offset}}$ accounts for the bias voltage. $\eta$ corresponds to the quantization step associated with the bit depth.
Combining Eq. 2 and Eq. 3, the ideal model for the optical-to-digital conversion is modeled as:
[…]
where $D$ signifies the pixel value in the raw image, $V_{\textrm{offset}}/{\eta}$ accommodates artificially introduced bias voltage to prevent output signals below 0. The raw response value of the bias voltage (i.e., black level) can be directly read out. When the dark current is 00, or we subtract the raw response value of the bias voltage, the pixel value in the raw image can be expressed as:
[…]
We observe from Eq. 5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly change the camera raw response value. This linear characteristic allows us to achieve an equivalent result to modifying the exposure time ($T$) by simulating the gain, thereby obtaining a set of bracketed exposure raw images. However, there is unavoidable noise in the actual imaging process. Therefore, we follow the existing denoising methods [39, 32] and try to avoid the effect of noise as much as possible during data acquisition (in Sec. 3.2.1).
3.2 Formation of Short- and Long-exposure Raw Pairs
Compared to sRGB images, HDR reconstruction from raw images has the advantages of more original information, simpler ISP processing, and less computation, making it a promising paradigm to deploy in edge devices. To this end, we construct a new raw HDR dataset with LDR-HDR data pairs, named RealRaw-HDR.
3.2.1 Data Acquisition
Based on the analysis in Sec. 3.1, we find that changing the sensor digital Gain, $K_{\textrm{d}}$, can achieve a similar luminance to adjusting the exposure time $K_{\textrm{a}}$ on the noise-free condition. Consequently, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images. The camera has 15.5 stops of dynamic range (15.5 bit) and a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 15 Pro Max primary camera single pixel size is only $1.22\mu m$). We also set the camera ISO to 800 or below and turned on the noise reduction feature to enhance image quality. At this point, the captured raw image has a low noise level.
@@ 3.2.1 Data Acquisition
Figure 1: The raw LDR-HDR pair formation pipeline. Two clean high dynamic range raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, and normalization. After manual digital gain, clip, add noise, and normalization, the long-exposure image $I_{l}$ is overexposed in the bright areas, and the short-exposure image $I_{s}$ dark area information is covered by noise.
3.2.2 Data Processing
@@
Figure 2: Real samples collected by the proposed raw LDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
Selection of exposure time ratio and initial adjustment. Commencing the pipeline, we pack two clean Bayer raw images. We select an exposure time ratio $r$ from {4, 8, 16} at this stage. The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, multiply by the maximum pixel value (Max: $2^{12}\times r$). This operation yields $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{2}$ correspondingly.
@@
Ground truth image ($I_{gt}$). HDR aims to recover detailed information from LDR images in both brighter and dark areas. Therefore, for a dual-exposure HDR sensor, we aim to recover the darkest areas of the short-exposure image from the long-exposure image. Thus, based on this principle, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$. The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions than $I_{s}$. As a result, our data formation pipeline efficiently generates an extensive array of LDR-HDR data pairs.
Luminance alignment. Finally, after the luminance alignment [14], we obtain the noisy raw LDR images $I^{{}^{\prime}}_{l}$ and $I^{{}^{\prime}}_{s}$, and the corresponding clean raw HDR image $I^{{}^{\prime}}_{gt}$.
@@
3.2.3 RealRaw-HDR Dataset
Our dataset is meticulously crafted for dual-exposure HDR sensors, supporting mainstream sensors, including Sony IMX327, IMX385, IMX585, and OV OS05B. To the best of our knowledge, there is an absence of a raw HDR dataset explicitly tailored for these HDR sensors. Our proposed data formation pipeline is efficient and user-friendly, enabling the creation of many high-quality data pairs effortlessly. We gather $240$ pairs of $8192\times 6192$ high-resolution raw image pairs and expand to 720 pairs. Fig. 2 shows an example of two generated LDR-HDR pairs with different exposure ratios. Additionally, by attaching an ISP pipeline to the end of our pipeline, we can create an sRGB-based HDR training dataset. In Tab. 1, we compare the statistics of our dataset with those of other existing HDR datasets. In this paper, all raw images have been processed with a fixed ISP, and HDR images are processed with the same tone mapping operator to obtain the sRGB version for visualization.
Table 1: The statistics comparison between Kalanatri [14], Chen [4] and our RealRaw-HDR dataset.

### v6

@@ Abstract
The growing prevalence of high-resolution displays on edge devices has created a pressing need for efficient high dynamic range (HDR) imaging algorithms. However, most existing HDR methods either struggle to deliver satisfactory visual quality or incur high computational and memory costs, limiting their applicability to high-resolution inputs (typically exceeding 12 megapixels). Furthermore, current HDR dataset collection approaches are often labor-intensive and inefficient. In this work, we explore a novel and practical solution for HDR reconstruction directly from raw sensor data, aiming to enhance both performance and deployability on mobile platforms. Our key insights are threefold: (1) we propose RepUNet, a lightweight and efficient HDR network leveraging structural re-parameterization for fast and robust inference; (2) we design a new computational raw HDR data formation pipeline and construct a new raw HDR dataset, RealRaw-HDR; (3) we design a plug-and-play motion alignment loss to suppress ghosting artifacts under constrained bandwidth conditions effectively. Our model contains fewer than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU. While being highly efficient, our model also achieves comparable performance to state-of-the-art HDR methods in terms of PSNR, SSIM, and a color difference metric.
I Introduction
@@
Most resource-constrained cameras exhibit standard dynamic range (SDR), rendering them unable to capture the full range of brightness and color information in real-world scenes. Conversely, high dynamic range (HDR) imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in SDR images. Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users. As a result, there has been an increasing focus on fusion-based HDR imaging methods.
@@
Recent methods [47, 24, 5, 31, 46] based on convolution neural networks (CNNs) [22, 51, 40, 16] have made impressive progress in HDR reconstruction performance, thanks to their scalability and flexibility from constructing elementary building blocks like convolutional layers. However, superior performance is usually obtained at a cost of a heavy computational burden [24, 46, 25]. Although this can be alleviated by elaborate network structures or dedicated computing engines (e.g., GPU and NPU), the hardware cost and power consumption still limit the deployment of existing deep HDR reconstruction networks. Specifically, the growing number of high-definition screens on edge devices (e.g., smartphones, security cameras, and televisions) calls for a practical HDR reconstruction solution.
@@
On the other hand, in the image processing pipeline, HDR reconstruction is widely used in the sRGB domain. Previous methods [47, 26, 48] exploit a set of sRGB images with different exposure levels to produce an HDR image, which has made rapid development in recent years. However, they tend to overlook three critical aspects. 1) Dataset Collection: Existing methods [4, 59] follow Kalantari et al. [18] to construct datasets. They first make the subject static and take three sets of images with different exposures, and then make the subject move twice to take dynamic images with different exposures. However, this process is labor-intensive and difficult to acquire on a large scale. 2) ISP Processing Speed: When obtaining raw SDR images with different exposures, the ISP pipeline [45, 9] must be performed separately on each exposure. This incurs additional memory and computational overhead and leads to lower frame rates for HDR image output. 3) Reconstruction Quality: Raw images contain more delicate details of the original sensor signal that can be lost while processing sRGB images. The limitations of current HDR reconstruction methods highlight the need for further research and development in HDR imaging.
@@
In this paper, we propose an efficient scheme for HDR reconstruction in the raw image domain. By analyzing the HDR image sensor system, we design a lightweight and efficient model for raw HDR reconstruction named RepUNet. RepUNet adopts reparameterization techniques and does not contain explicit computationally expensive alignment modules, such as optical flow [18], deformation convolution [5], or attention [24, 49, 39], which are commonly used in existing deep learning-based HDR reconstruction methods [18, 24, 44, 30]. To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss, which encourages the network to focus on local motion patterns and alleviate misalignment between short- and long-exposure images. Consequently, our approach significantly reduces hardware costs and improves the real-time performance of HDR imaging systems.
@@
To further promote the HDR imaging system, we investigate the HDR sensor imaging principle. We observe that changing the Gain of the image sensor can have a similar effect as modifying the exposure time under noise-free conditions. Leveraging this insight, we design an automatic control imaging system that captures raw images with different exposures, based on a digital camera photoelectric signal conversion model. This automatic control operable system satisfies real-world scenes’ dynamic range requirements, making it a practical tool for generating high-quality HDR images. The resulting RealRaw-HDR dataset includes many SDR-HDR pairs for training and evaluation. By incorporating the unique characteristics of raw images into our approach, we can achieve superior HDR reconstruction results with increased efficiency and accuracy.
@@
    We investigate the structure reparameterizable technique for the HDR task and propose a lightweight model, RepUNet, with Topological Convolution Block (TCB). TCB can be used to improve the HDR performance of any HDR model without introducing any extra burden for inference.
@@
Our contributions represent a significant step forward in raw HDR image reconstruction research, providing an effective and efficient solution for producing high-quality HDR images. Our model contains fewer than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU. While highly efficient, our model also outperforms the state-of-the-art HDR methods by a large margin in terms of PSNR, SSIM, and a color difference metric.
II Related Work
II-A HDR Imaging
@@
Recently, benefiting from the fast development of deep learning techniques, training deep neural networks for effective HDR reconstruction has become increasingly popular. Many methods apply deep neural networks [20, 30, 35, 17] to learn the production of high-quality HDR images from a set of bracketed exposure SDR images. Kalantari et al. [18] proposed a CNN-based HDR approach that employs optical flow to align SDR sRGB images before network inference. Wu et al. [44] approached HDR imaging as an image translation problem without explicit motion alignment. Yan et al. [46] introduced spatial attention to achieve SDR image alignment. Liu et al. [24] presented an attention-guided deformable convolutional network for multi-frame HDR imaging. Niu et al. [31] proposed a multi-frame HDR imaging method based on generative adversarial learning. Liu et al. [25] proposed a Transformer-based [32] HDR imaging method. SCTNet [37] proposed an alignment-free architecture employing semantic consistency transformers. These deep learning-based approaches [24, 25] have consistently pushed the boundaries of state-of-the-art performance. To address the need for deployment on mobile devices, Prabhakar et al. [33] introduced an efficient method for generating HDR images using a bilateral guided up-sampler and exploring zero-learning for HDR reconstruction. CEN-HDR introduced a lightweight architecture designed for real-time mobile applications. However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset. They overlook the large computational and storage resources the ISP pipeline requires to process bracketed exposure raw images. It also complicates the ISP system, making it challenging for resource-limited cameras to output high-quality video/images.
@@
To simplify the ISP system, another class of HDR reconstruction methods is based on raw image input. Google HDR+ [11] produced the raw HDR image by aligning and merging a burst of raw frames with the same low exposure. Nevertheless, this approach requires a complex ISP system design and takes up a lot of DDR memory. Zou et al. [59] proposed reconstructing HDR images from a single raw image and collecting a raw/HDR paired dataset. However, this dataset is not suitable for real HDR sensors. Therefore, none of the existing methods can meet the requirements of real scenarios.
II-B Low-level Raw Image Processing
@@
Due to the merits of raw data, raw-based image processing [21, 53] has made significant progress in recent years. The work in [54] first performs the demoiréing task in the raw domain and then utilizes a pre-trained ISP module to transform the result into the sRGB domain. Yang et al. [50] proposed a single-stage network empowered by feature domain adaptation to decouple the denoising and color mapping tasks in raw low-light enhancement. Zhang et al. [56] constructed a real-world super-resolution dataset by designing an optical zoom system and proposed a baseline network with a bilateral contextual loss. Qian et al. [34] solved the joint demosaicing, denoising, and super-resolution task with the raw input. Wang et al. [41] proposed a lightweight and efficient network for raw image denoising. Sharif et al. [1] proposed a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors. Wei et al. [43] investigated the low-light image denoising considering the image sensor photoelectric properties. Yue et al. [52] achieved state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs. Zou et al. [60] proposed a model tailor-made for Raw images, harnessing the unique features of Raw data to facilitate the Raw-to-HDR mapping. Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data. However, acquiring paired data in the raw domain is difficult and expensive. Our work proposes a new large-scale, high-quality raw dataset and provides a data synthesis pipeline to acquire raw SDR-HDR pairs based on the HDR sensor imaging system.
III New Dataset Formation Pipeline
@@
III-A Analysis of CMOS Imaging System
@@
Subsequently, the camera amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [38, 58]:
@@
where $K_{\textrm{a}}$ represents the Analog Gain, $K_{\textrm{d}}$ stands for the Digital Gain, and $V_{\textrm{offset}}$ accounts for the bias voltage. $\eta$ corresponds to the quantization step associated with the bit depth. Combining Eq. 2 and Eq. 3, the ideal model for the optical-to-digital conversion is modeled as:
$D=\frac{K_{\textrm{a}}T\overline{S}A\int_{\lambda}E(\lambda)q(\lambda)d\lambda+V_{\textrm{offset}}}{\eta}\times K_{\textrm{d}},$ (4)
where $D$ signifies the pixel value in the raw image, $V_{\textrm{offset}}/{\eta}$ accommodates artificially introduced bias voltage to prevent output signals below 0. The raw response value of the bias voltage (i.e., black level) can be directly read out. When the dark current is 0, or we subtract the raw response value of the bias voltage, the pixel value can be expressed as:
$D=\frac{K_{\textrm{a}}T\overline{S}A\int_{\lambda}E(\lambda)q(\lambda)d\lambda}{\eta}\times K_{\textrm{d}}.$ (5)
We observe from Eq. 5 that under noise-free conditions, adjusting the gain ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly change the camera raw response value. This linear characteristic allows us to achieve an equivalent result to modifying the exposure time ($T$) by simulating the gain, thereby obtaining a set of bracketed exposure raw images. However, there is unavoidable noise in the actual imaging process. Therefore, we follow the existing denoising methods [52, 42] and try to avoid the effect of noise as much as possible during data acquisition.
III-B Formation of Short- and Long-exposure Raw Pairs
Compared to sRGB images, HDR reconstruction from raw images has the advantages of more original information, simpler ISP processing, and less computation, making it a promising paradigm to deploy in edge devices. To this end, we construct a new raw HDR dataset with SDR-HDR data pairs, named RealRaw-HDR.
III-B1 Data Acquisition
Based on the analysis in Sec. III-A, we find that changing the sensor digital Gain, $K_{\textrm{d}}$, can achieve a similar luminance to adjusting the exposure time $K_{\textrm{a}}$ on the noise-free condition. Consequently, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images. The camera has 15.5 stops of dynamic range (15.5 bit) and a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 15 Pro Max primary camera single pixel size is only $1.22\mu m$). We also set the camera ISO to 800 or below and turned on the noise reduction feature to enhance image quality. At this point, the captured raw image has a low noise level.
@@ III-B1 Data Acquisition
Figure 1: The raw SDR-HDR pair formation pipeline. Two clean HDR raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction and normalization. After manual digital gain, clip, add noise, and normalization, the long-exposure image $I_{l}$ is overexposed in the bright areas, and the short-exposure image $I_{s}$ has dark area information covered by noise.
III-B2 Data Processing
@@
Figure 2: Real samples collected by the proposed raw SDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
Selection of exposure time ratio and initial adjustment. Commencing the pipeline, we pack two clean Bayer raw images. We select an exposure time ratio $r$ from {4, 8, 16} at this stage. The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, are multiplied by the maximum pixel value (Max: $2^{12}\times r$). This operation yields $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{2}$ correspondingly.
@@
Ground truth image ($I_{gt}$). HDR aims to recover detailed information from SDR images in both brighter and darker areas. Therefore, for a dual-exposure HDR sensor, we aim to recover the darkest areas of the short-exposure image from the long-exposure image. Thus, based on this principle, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$. The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions than $I_{s}$. As a result, our data formation pipeline efficiently generates an extensive array of SDR-HDR data pairs.
Luminance alignment. Finally, after the luminance alignment [18], we obtain the noisy raw SDR images $I^{{}^{\prime}}_{l}$ and $I^{{}^{\prime}}_{s}$, and the corresponding clean raw HDR image $I^{{}^{\prime}}_{gt}$.
@@
III-C RealRaw-HDR Dataset
Our dataset is meticulously crafted for dual-exposure HDR sensors, supporting mainstream sensors, including Sony IMX327, IMX385, IMX585, and OV OS05B. To the best of our knowledge, there is an absence of a raw HDR dataset explicitly tailored for these HDR sensors. Our proposed data formation pipeline is efficient and user-friendly, enabling the creation of many high-quality data pairs effortlessly. We gather $240$ pairs of $8192\times 6192$ high-resolution raw image pairs and expand to 720 pairs. Fig. 2 shows an example of two generated SDR-HDR pairs with different exposure ratios. Additionally, by attaching an ISP pipeline to the end of our pipeline, we can create an sRGB-based HDR training dataset. In Tab. I, we compare the statistics of our dataset with those of other existing HDR datasets. In this paper, all raw images have been processed with a fixed ISP, and HDR images are processed with the same tone mapping operator to obtain the sRGB version for visualization.
TABLE I: The statistics comparison between Kalanatri [18], Chen [5] and our RealRaw-HDR dataset.